a reading center demands that the students will perform better on a standardized reading test after going through the reading test after going through the reading course offered by their center. the table shows the reading scores of 5 students before and after the course. at α=0.10, is there enough evidence to conclude that the students’ scores after the course are better than the scores before the course?


STUDENT BEFORE AFTER

1 85 88

2 96 85

3 70 89

4 76 86

5 81 92



1
Expert's answer
2022-05-24T17:37:32-0400

"H_0: \mu_d\le0\\
H_a:\mu_d>0" (Claim)

d=(score before-score after)




"\=d=\frac{\sum{d}}{n}=-32/5=-6.4"

"s_d=\sqrt{\frac{n(\sum{d^2})-(\sum{d})^2}{n(n-1)}}=\sqrt{\frac{5(712)-1024}{20}}=11.26"

We can find critical value "t_0" using t-table (in this problem "\alpha=0.10, df=5-1=4" ):


"t_0=1.533"

The standardized test statistic is:

"t=\frac{\=d-\mu_d}{s_d/\sqrt{n}}=\frac{-6.4-0}{11.26/\sqrt5}=-0.254"

Since "t<t_0 (-0.254<1.533)", "H_0" is not rejected.

So there is not enough evidence at the 10% level to support the claim that the student`s scores after the course are better than the scores before the course.